cooper and nickel can be mixed to form a corrosion resistant alloy. if a particular alloy is prepared by mixing 13.7 moles of cooper with 1.00 moles of nickel. what is the mass percentage of nickel in the alloy
Answer #1

Answer :

Molar mass of Cu = 63.5 g /mol

1mol of Cu has mass = 63.5 g

13.7 mole of Cu has mass = 13.7 * 63.5 = 869.95 g

Molar mass of Ni = 58.69 g/mol

1 mol of Ni has mass = 58.69 g

Total mass of alloy = mass of 13.7 moles of Cu + mass of 1 mol of Ni = 869.95 + 58.69 = 928.64 g

% of Ni in the alloy = (mass of Ni / total mass of alloy) * 100 = (58.69 / 928.64 ) * 100 = 6.32 %